OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x @bibby @smexygirl97 @adrynicoleb @tester97 @Darrius

OpenStudy (anonymous):

@jojo4eva @Jonathan1997 @BOOKER23

OpenStudy (anonymous):

@dinmosiaren

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

@Jimmysixth

OpenStudy (anonymous):

Do you mean: \[\frac{ 4 }{ 5x } + \frac{ 4 }{ 3 } = 2x\] or \[\frac{ 4 }{ 5 }x + \frac{ 4 }{ 3 } = 2x\] ?

OpenStudy (anonymous):

The second is easy: \[\frac{ 4 }{ 5 }x + \frac{ 4 }{ 3 } = 2x\] \[15*\frac{ 4 }{ 5 }x +15* \frac{ 4 }{ 3 } = 15*2x\] \[12x+20=30x\] \[20=30x-12x\] \[20=18x\] \[x=\frac{ 20 }{ 18 }=\frac{ 10 }{ 9 }=1\frac{ 1 }{ 9 }\]